The human ear can detect sounds with frequencies between 20 Hz and 16 kHz. Find the
greatest and the smallest wavelengths the ear can detect, assuming the sound travels.
1
Expert's answer
2011-08-25T13:35:35-0400
The LENGTH of the WAVE is defined as distance between two next points of a sound wave which are in identical oscillatory position (have an identical phase).
The interrelation between length of a wave and frequency is defined by the following formula: λ = c / f , where "c" - speed of a sound in air, is usually used value of 330 (340) m/s. (The given value depends on temperature and humidity of air, and also from barometric pressure), and "f" - frequency in Hz.



16 kHz=16000Hz
The smallest wavelength λsmall = 330/16000 = 0.50625 m.
The greatest wavelength λgreat = 330/20 = 16.5 m.
The smallest wavelength is 0.020625 m.
The greatest wavelength is 16.5 m.